0 3.2.69) Find the x-intercepts and y-intercepts of the circle x2 + y2-4x- 6y +4 3.4.25) Find the domain of f (x)9x 3.6.1) Find the equation of any parabola that has vertex V(-3, 1) x+1
Answer #1

3.2.69)
The given equation is:
x^2 + y^2 - 4x - 6y + 4 = 0

to get y intercepts, put x=0
0^2 + y^2 - 4*0 - 6y + 4 = 0
y^2 - 6y + 4 = 0

This is quadratic equation (ay^2+by+c=0)
a = 1
b = -6
c = 4

Roots can be found by
y = {-b + sqrt(b^2-4*a*c)}/2a
y = {-b - sqrt(b^2-4*a*c)}/2a

b^2-4*a*c = 20

roots are :
y = 5.236 and y = 0.7639

These are y intercepts
Answer: y intercepts are:
y = 5.2 and y = 0.76

b)
To get x intercepts, put y=0
x^2 + y^2 - 4x - 6y + 4 = 0
x^2 + 0^2 - 4x - 6*0 + 4 = 0
x^2 - 4x + 4 = 0
(x-2)^2 = 0
x = 2
This is x intercepts
Answer: x intercept is:
x=2

I am allowed to answer only 1 question at a time